# Executable Knowledge Graphs for Replicating AI Research

### v1

@@ 3.1 Settings
We evaluate xKG on the lite collection of PaperBench Code-Dev using a structured rubric Starace et al. (2025), a weighted tree of binary criteria whose leaves are aggregated by an o3-mini-based evaluator into a single score. We integrate xKG into BasicAgent (a ReAct-style agent), IterativeAgent (adds a self-improvement loop), both with a one-hour runtime limit, and PaperCoder(a repository-level reproduction agent with a fixed workflow). See Appendix A for more details.
@@ 3.2 Main Results
As shown in Table 1, xKG achieves substantial performance gains across diverse agent frameworks and LLM backbones. On the general ReAct-style IterativeAgent with DeepSeek-R1, xKG delivers a performance improvement of 8.20%. The effectiveness of xKG is further highlighted by the 10.90% improvement achieved with PaperCoder powered with o3-mini. Notably, the impact of xKG is also highly paper-dependent. While BasicAgent with o3-mini achieves a remarkable 24.26% performance gain on MU-DPO, the same configuration yields only a 2.58% improvement on One-SBI and even a 0.15% drop on the FRE task. This striking contrast reveals a critical dependency on the target paper (details in Appendix B).
@@ 3.3 Further Analysis
Table 2: Ablation study on the core components of xKG. Performance is averaged over all 5 papers.
@@
As shown in Table 2, our ablation study conducted on PaperCoder framework with o3-mini setup, reveals that removing any component degrades performance. The most significant drop occurs when removing Code Nodes, decreasing the score by 4.56% (53.21% $\rightarrow$ 48.65%), suggesting that LLM agents benefit immensely from fine-grained knowledge, with executable code being the most critical component. Ablating Paper Nodes yields a substantial degradation of 2.13%, highlighting the value of a high-level structural overview of the target task. In contrast, omitting Technique Nodes results in a modest 1.05% drop, since the function of each technique is already implicitly captured by the Code Nodes, rendering the explicit description redundant.
@@
Building on the above findings, we conduct a further analysis into how the quality of Code Nodes within xKG influences performance. Using PaperCoder with o3-mini on two high-gain papers, MU-DPO and TTA-FP, we compare xKG with four configurations, each repeated three times to mitigate stochasticity (Figure 2): w/o Code, without access to any code nodes; + Raw Code, which incorporates code nodes with raw retrieved snippets; + Rewrite, using LLM-rewritten executable nodes but omitting the verification step.
@@
As illustrated in Figure 2, our full approach not only achieves the highest mean score but also exhibits low variance. Notably, even incorporating raw code snippets (+ Raw Code) improves performance, validating that our method effectively localizes necessary code. A critical insight emerges from the + Rewrite setting, which underperforms even the raw snippet baseline. We attribute this to a misleading guidance phenomenon: well-formatted but contextually irrelevant code can cause the agent to blindly adopt the retrieved snippets, deviating from the target paper’s specific implementation.
@@
To understand the mechanism behind the performance gains, we conduct a case study on the MU-DPO paper (Figure 3). We notice that xKG enriches information granularity, allowing agents to generate critical details accurately, and improves modular implementation capability, enabling agents to reuse verified code for functionally correct implementations, as illustrated by the case colors in Figure 3.
Figure 3: Case Study on MU-DPO. A comparison of performance with and without xKG on IterativeAgent.
@@ 4 Conclusion
We introduce xKG, which improves AI research replication. We aim for xKG to serve as an AI-for-Research knowledge base, reducing noise from web retrieval and improving efficiency.
@@
Appendix A Experimental Setup
@@ A.1 Benchmarks
Table 3: Abbreviations and statistics for the PaperBench tasks evaluated in this work. We assign a brief abbreviation to each paper for easier reference. The "CodeDev Nodes" column specifies the number of nodes to evaluation for each reproduction task.
The original PaperBench benchmark Starace et al. (2025) is designed to evaluate the ability of AI agents to reproduce state-of-the-art AI research from scratch. The full benchmark includes 20 recent papers from top-tier machine learning conferences (e.g., ICML 2024), where agents must understand each paper, develop a complete codebase, and replicate its empirical results.
@@
In our study, we adopt the pre-defined lite subset of PaperBench Code-Dev provided in the official repository, spanning diverse AI domains including machine learning (ML), reinforcement learning (RL), and natural language processing (NLP). Furthermore, since PaperBench shows that BasicAgent and IterativeAgent gain little performance improvement after one hour, we cap their execution time at one hour for efficiency and cost reasons.
@@
A.2 Configuration Details
The configuration of our xKG framework comprises both hyperparameters and prompts. The hyperparameters are managed via a central config.yaml file, which is organized into modules for Code-RAG, Paper-RAG, and Knowledge Graph Retrieval. We summarize the key parameters for each module in Tables 4-7. In addition, the specific prompts designed in our system are detailed in Appendix C.
Appendix B Further Analysis on Target Paper
@@
More fundamentally, this performance disparity is tied to the paper’s research archetype. analytical papers, such as MU-DPOLee et al. (2024), which synthesize and refine existing techniques, benefit substantially as their components are well-represented in xKG. Conversely, methodological papers like One-SBIGlöckler et al. (2024), which introduce fundamentally novel architectures, find less directly applicable knowledge, as their core innovations have limited precedent in the corpus. This outcome is logical, as the performance bottleneck shifts from knowledge argumentation to the intrinsic innovative capacity of the base LLM itself.
@@
Appendix C Prompts
@@
C.1 Paper Parsing Prompts
@@
C.2 Code Repository Parsing Prompts
@@
C.3 Knowledge Graph Construction Prompts
@@
Appendix D Running Examples of xKG
Figure 5: An example of xKG data storage. Paper Nodes are stored as JSON files, with technique and code nodes embedded as structured dictionaries, where key-value pairs are used to create a one-to-one mapping representing the implementation relationship.

### v2

@@ 3.1 Settings
We evaluate xKG on the lite collection of PaperBench Code-Dev Starace et al. (2025), a benchmark for repository-level paper reproduction from scratch, scored by an o3-mini-based evaluator using a weighted, tree-structured rubric. We integrate xKG into BasicAgent (a ReAct-style agent), IterativeAgent (adds a self-improvement loop), both with a one-hour runtime limit, and PaperCoder(a powerful agent tailored for repository-level paper reproduction.). See Appendix A for more details.
@@ 3.2 Main Results
As shown in Table 1, xKG achieves substantial performance gains across diverse agent frameworks and LLM backbones. On the general ReAct-style IterativeAgent with DeepSeek-R1, xKG delivers a performance improvement of $\uparrow$8.20%. The effectiveness of xKG is further highlighted by the $\uparrow$10.90% improvement achieved with PaperCoder powered with o3-mini, underscoring its broad applicability from simpler agents to more advanced ones. Notably, the impact of xKG is also highly paper-dependent. While BasicAgent with o3-mini achieves a remarkable 24.26% performance gain on MU-DPO, the same configuration yields only a 2.58% improvement on One-SBI and even a 0.15% drop on the FRE task, revealing a critical dependency on the target paper (details in Appendix C).
@@ 3.3 Further Analysis
Table 2: Ablation study on xKG node types.
@@
As shown in Table 2, our ablation study conducted on PaperCoder with o3-mini reveals that removing any component degrades performance. The most significant drop occurs when removing Code Nodes, decreasing the score by 4.56% (53.21% $\rightarrow$ 48.65%), suggesting that LLM agents benefit immensely from fine-grained knowledge, with executable code being the most critical component. Ablating Paper Nodes yields a substantial degradation of 2.13%, highlighting the value of a structural overview of the target paper. In contrast, omitting Technique Nodes results in a modest 1.05% drop, since the function of each technique is already captured by the Code Nodes, rendering the explicit description redundant.
@@
Building on the above, we conduct a further analysis of how Code Nodes influence performance. We analyze the impact of code quality on PaperCoder (o3-mini), comparing four settings with 3 runs each: w/o Code (no code), + Raw Code (raw snippets), + Rewrite (LLM-rewritten but unverified), and + Verify (LLM-rewritten & verified).
@@
Techs. Codes Tech-Code Pairs
Valid Rate (%) 89.44 100.00 74.51
Table 3: Human evaluation of xKG quality.
Figure 3: Comparison of invalid Technique Nodes in xKG before and after knowledge filtering.
As illustrated in Figure 2, our full method excels in score and stability. Notably, even incorporating raw code snippets (+ Raw Code) significantly improves performance, validating that our method effectively localizes necessary code. The failure of + Rewrite ablation reveals a key insight: agents are misled by well-formatted, semantically similar yet technically irrelevant knowledge, a phenomenon consistent with previous findings Wu et al. (2024). Our LLM Verifier mitigates this by prioritizing technical relevance over semantic similarity, filtering out distracting irrelevant information, thus boosting the score improvement varying from 5.75 to 8.20 percentage points.
Automated Construction architects robust xKG.
Human evaluation (Table 3) confirms the high quality of xKG with 89.44% of Techniques representing self-contained academic concepts, all Code Nodes executable, and 74.51% of Tech-Code pairs exactly matched. During xKG construction, the Knowledge Filtering step mitigates LLM errors (e.g., overspecificity, hallucinations) by pruning techniques with no code retrieved (Figure 3). Concurrently, a self-debugging loop achieves 100% code executability, up from an initial 52.38%. Tech-Code mismatches primarily stem from broader code snippets with little impact on application. See Appendix B for more details of constructed xKG.
@@
To understand the mechanism behind the performance gains, we conduct a case study on the MU-DPO (Figure 4). This case study reveals that xKG transforms agents from dummy scaffolders to substantive implementers, arming them with both the precise method details for accurate planning and the verified, modular reference code for robust implementation (cf. Figure 4).
Figure 4: Case Study on MU-DPO. Highlights illustrating agent implementation contrast with/without xKG.
@@ 4 Conclusion
We introduce Executable Knowledge Graphs (xKG) to make implicit research knowledge modular and executable, boosting agent replication performance. Looking forward, we envision xKG serving not only as a dynamic knowledge base, but also as a flexible instrument to accelerate AI Research with improved efficiency and verifiability.
@@
Appendix A Experimental Details
@@ A.1 Benchmarks
Table 4: Abbreviations for the PaperBench tasks evaluated in this work.
Stage LLM API Calls (avg) Input Tokens (avg) Output Tokens (avg) Cost (avg, \$)
Technique Extraction o4-mini 8.21 53,142.36 11,523.43 0.1092
Code Modularization o4-mini 33.36 202,625.21 91,426.14 0.6252
Knowledge Filtering
Total o4-mini 41.57 255,767.57 102,949.57 0.7344
Table 5: Average cost analysis for constructing the xKG for a single paper.
The original PaperBench benchmark Starace et al. (2025), featuring publicly available tasks and evaluation rubrics, is designed to evaluate the ability of AI agents to reproduce AI research from scratch.
@@
In our study, we adopt the pre-defined lite subset of PaperBench Code-Dev provided in the official repository, spanning diverse AI domains including machine learning, reinforcement learning, and natural language processing. Furthermore, since PaperBench shows that BasicAgent and IterativeAgent gain little performance improvement after one hour, we cap their execution time for efficiency and cost.
@@
A.2 Configuration
The configuration of our xKG framework comprises both hyperparameters and prompts. The hyperparameters are managed via a central config.yaml file, which is organized into modules for Code-RAG, Paper-RAG, and Knowledge Graph Retrieval. We summarize the key parameters for each module in Tables 9-12. In addition, the specific prompts designed in our system are detailed in Appendix E.
A.3 Cost Evaluation
Our xKG construction pipeline leverages OpenAI’s o4-mini via third-party API, chosen for its strong cost-effectiveness and robust text/code capabilities. We quantify the cost of this process in Table 5.
As shown, for an average cost of about \$0.73 per paper, our pipeline transforms a scientific publication into an executable knowledge graph. The most significant cost driver is the iterative self-debugging loop of the Code Modularization stage, simultaneously serving as the critical quality-assurance mechanism for our code nodes. The key advantage of xKG lies in this one-time investment: each processed paper is converted into a durable and reusable knowledge resource.
Appendix B Analysis of constructed xKG
B.1 Self-Evolution of xKG
Domain Subfield Count
Generative AI Diffusion Models, Distribution Modeling, Controllable Generation 14
AI Safety & Robustness Test-Time Adaptation, Continual Learning, Toxicity Detection 13
Simulation-Based Inference Inverse Inference, Bayesian Inference, Scientific Computing 8
Reinforcement Learning Exploration Strategies, Agent Decision-making, Policy Optimization 7
Mechanistic Interpretability Circuit Analysis, Feature Visualization, Geometric Analysis 7
Table 6: Domains of the original 42 papers in xKG. The sum of counts exceeds the total number of papers due to interdisciplinary classifications.
Domain Subfield Count
Generative AI Diffusion Models, Distribution Modeling, Controllable Generation 18
AI Safety & Robustness Test-Time Adaptation, Continual Learning, Toxicity Detection 15
Simulation-Based Inference Inverse Inference, Bayesian Inference, Scientific Computing 8
Mechanistic Interpretability Circuit Analysis, Feature Visualization, Geometric Analysis 8
Reinforcement Learning Exploration Strategies, Agent Decision-making, Policy Optimization 7
Model Adaptation & Efficiency Transfer Learning, Visual Prompting, Model Reprogramming 4
Table 7: Updated domains of the 56 papers in the expanded xKG, where some papers span multiple domains.
xKG is not a static knowledge base but a flexible framework that automates the collection, extraction, and validation of knowledge in a paper-aware way. To illustrate this, we select two additional target papers from PaperBench, bridging-data-gaps and sample-specific-masks, to demonstrate the dynamic evolution of xKG.
The expansion from a 42-paper corpus (Table 6) to a 56-paper version (Table 7), curated from publicly available research, highlights the autonomous evolution of xKG. When new target papers are introduced, our Corpus Curation module automatically gathers relevant literature and completes the full construction and verification cycle. This adaptability substantiates our claim that xKG is an adaptive system, effectively handling scenarios like fetching updated research or leveraging broader domain knowledge.
Paper HumanKG Nodes xKG Nodes Precision (w) Recall (w) F1-Score (w)
CFDG 4 3 1.00 0.90 0.95
RND 11 7 0.92 0.86 0.89
TENT 4 4 1.00 1.00 1.00
Table 8: Comparison of automatically-constructed xKG and human-constructed KG.
B.2 Comparison to Human-Constructed KG
To further validate the quality of our automated xKG paper deconstruction, we manually construct KGs for several papers and compared them against the nodes generated by xKG. Recognizing that manual annotation is a time-consuming process requiring significant domain expertise, we randomly selected three papers from the xKG corpus and invited expert PhD candidates to annotate them.
As shown in Table 8, the high weighted F1-scores demonstrate that our automated KG construction closely aligns with human annotation. Crucially, manually annotating a single paper takes 30-60 minutes, which highlights the significant efficiency gains of our automated pipeline.
B.3 Analysis of Paper Represent Granularity
Our hierarchical graph design (Section 2.1) is intentionally flexible, built to preserve a paper’s natural methodological structure rather than forcing a uniform method decomposition. While most papers (71%) in xKG are decomposed into 2-5 techniques, a significant portion of papers are represented as single, atomic nodes (12%) or complex works with 6+ nodes (17%), ensuring that potentially useful information from high-level methodologies to fine-grained details is captured at its appropriate level.
Appendix C Analysis on Target Paper
@@
More fundamentally, this performance disparity is tied to the paper’s research archetype. analytical papers, such as MU-DPOLee et al. (2024), which synthesize and refine existing techniques, benefit substantially as their components are well-represented in xKG. Conversely, methodological papers like One-SBIGlöckler et al. (2024), which introduce fundamentally novel architectures, find less directly applicable knowledge, as their core innovations have limited precedent in the corpus. This outcome is logical, as the performance bottleneck shifts from knowledge augmentation to the intrinsic innovative capacity of the base LLM itself.
@@
Appendix D Running Examples of xKG
Figure 6: An example of structural xKG data storage. Paper Nodes are stored as JSON files, with technique and code nodes embedded as structured dictionaries, where key-value pairs are used to create a one-to-one mapping representing the implementation relationship.
Appendix E Prompts
@@
E.1 Paper Parsing Prompts
@@
E.2 Code Repository Parsing Prompts
@@
E.3 Knowledge Graph Construction Prompts